OpenStudy (anonymous):

If sinθ = 3/4 what are the values of cos θ. Use: sin^2 θ + cos^2 θ = 1

OpenStudy (anonymous):

@SithsAndGiggles @e.mccormick @texaschic101 @Luigi0210 @zepdrix @nincompoop Help please :)

OpenStudy (anonymous):

@triciaal @chucho78 @wrstlr3232 I really need help with this.

zepdrix (zepdrix):

\[\Large\rm \color{orangered}{\sin \theta=\frac{3}{4}}\]Plugging this into our Pythagorean Identity,\[\Large\rm \color{orangered}{\sin}^2\color{orangered}{\theta}+\cos^2\theta=1\]gives us,\[\Large\rm \left(\color{orangered}{\frac{3}{4}}\right)^2+\cos^2\theta=1\]

zepdrix (zepdrix):

And then we need to solve for cosine from there. Understand the plugging in and stuff? :o

OpenStudy (sachin00007):

take the square of both the sides. the sin and the 3/4. then solve for cos^2 by 1-(3/4)^2

OpenStudy (anonymous):

@zepdrix What do I do next?

OpenStudy (sachin00007):

and then take the square root of both the sides

OpenStudy (triciaal):